# 2-Cats: 2D Copula Approximating Transforms

### v3

@@ 5 Experimental Results
We now present our experimental results on different fronts: (1) validating our Empirical estimates for the first derivative, a crucial step when training 2-Cats; (2) evaluating 2-Cats on synthetic datasets; (3) and evaluations on real datasets.
Before presenting results, we note that our models were implemented using Jax [^3] https://jax.readthedocs.io/en/latest//Flax [^4] https://flax.readthedocs.io. Given that Jax does not implement CDF methods for Multivariate Normal Distributions, we ported a fast approximation of bivariateCDFs [Tsay and Ke, 2023] and employed it for our models that rely on Gaussian CDFs/Copulas. Our code is available at https://anonymous.4open.science/r/2cats-E765/. Baseline methods (below) were evaluated using the author-supplied source code for each method (details are on Appendix E).
@@ 5.1 Empirical First Derivative Estimator
We validate our empirical estimations for the first derivative of Copulas, namely: $\frac{\widehat{\partial C^{n}}(u,v)}{\partial u}$ and $\frac{\widehat{\partial C^{n}}(u,v)}{\partial v}$. To do so, we use the closed-form equations for the first derivatives of Gaussian, Clayton, and Frank Copulas described in  [Schepsmeier and Stöber, 2014].
Figure 2: First derivative estimation
These Copulas have a single parameter, which we varied. Gaussian Copulas ($\rho$) has a correlation parameter, and we set it to: 0.1, 0.5, and 0.9. Clayton and Frank copulas’ mixing parameter ($\theta$) was set to 1, 5, and 10. For each of these configurations, we measured the coefficient of determination ($R2$) between empirical estimates and exact formulae. One example of such an estimate for a Frank Copula with $\theta=10$ is shown in Figure 2. Overall, we found that R2 was above $0.899$ in every setting, validating our approach.
5.2 2-Cats on Synthetic Data
We now turn to our results on synthetic datasets. Our primary focus, as is the case on baseline methods, will be on capturing the data likelihood:
$f(x_{1},x_{2})=f(x_{1})\,f(x_{2})\,c(F^{-1}(x_{1}),F^{-1}(x_{2}))$
For both 2-Cats models, we estimate this density as:
$f(x_{1},x_{2})\approx k_{1}(x_{1})\,k_{2}(x_{2})\,\frac{\partial^{2}H_{\bm{\theta}}(u,v)}{\partial u\partial v}$
Here, $k_{1}(x_{1})$ and $k_{2}(x_{2})$ are KDE models for the PDFs for $X_{1}$ and $X_{2}$ respectivelly. The bandwidth for these methods is determined using Silverman’s rule [Silverman, 1986].
As is commonly done, we evaluate the natural log of this quantity as the data log-likelihood and use its negative to compare 2-Cats with other methods. With this metric of interest in place, we considered two 2-Cats models:
  1. 1.
    2-Cats-G (Final Layer Gaussian): 2-Cats with a Gaussian bivariateCDF as $G$.
  2. 2.
    2-Cats-L (Final Layer Logistic): 2-Cats with a Logistic bivariateCDF as $G$.
Thus, we considered two forms of $G$. The first was the CDF of a bivariate Gaussian Distribution [Tsay and Ke, 2023]. The second one is the CDF of the Flexible bivariate Logistic (see Section 11.8 of Arnold [1992]):
$F(x_{1},x_{2})=\Pr[X_{1}<x_{1},X_{2}<x_{2}]=$
$\big{(}(1+e^{-\alpha\frac{x_{1}-\mu_{1}}{\sigma_{1}}}+e^{-\alpha\frac{x_{2}-\mu_{2}}{\sigma_{2}}}+e^{-\alpha(\frac{x_{1}-\mu_{1}}{\sigma_{1}}+\frac{x_{2}-\mu_{2}}{\sigma_{2}})})^{\frac{1}{\alpha}}\big{)}^{-1}$
@@
Recall that the data density is proportional to the pseudo-likelihood. We primarily emphasized the likelihood loss ($w^{c}$). This accentuates our pivotal component. Notably, the scale of its values differs from that of the MSE components, necessitating a proportional adjustment in weight selection. In light of these factors, we fixed $w^{C}=0.01$, $w^{dC}=0.5$, and $w^{c}=0.1$. We note that these hyper-parameters were sufficient to show how our approach is better than the literature. Thus, no exhaustive grid search was needed.
Our models were trained using early stopping. Here, for each training epoch, we evaluated the pseudo-log-likelihood on the training set and kept the weights of the epoch with the best likelihood on the training dataset.
Table 1: Results for Synthetic Data. Copula parameters are shown on the header. Marginals come from two Normal distributions ($\mu=0,\sigma=1$). $95\%$ confidence intervals (bootstrap) show. Green indicates that a 2-Cats model was, on average, better than the best baseline (indicated by Red). Underscores highlight the best statistical ties with the best model.
@@
In both our synthetic and real data experiments, we considered the following approaches as Deep Learning baselines: Deep Archimedian Copulas (ACNET) [Ling et al., 2020]; Generative Archimedian Copulas (GEN-AC) [Ng et al., 2021]; Neural Likelihoods (NL) [Chilinski and Silva, 2020]; and Implicit Generative Copulas (IGC) [Janke et al., 2021]. ACNET, GEN-AC, and IGC are all methods focused on Copulas. ACNET, GEN-AC, and IGC present experiments for bivariate data on their studies. Here, we use the same hyperparameters employed by the authors. This is also true for NL, but the authors did not present an analysis of synthetic Copula datasets or bivariate data in their study. We consider NL a baseline given that it presents a similar idea to ours: employ monotonic NNs.
@@
We also considered several Non-Deep Learning baselines. These were the Parametric and Non-Parametric Copula Models from Nagler et al. [2017]. They are referenced as: The best Parametric approach from several known parametric copulas (Par), non-penalized Bernstein estimator (Bern), penalized Bernstein estimator (PBern), penalized linear B-spline estimator (PSPL1), penalized quadratic B-spline estimator (PSPL2), transformation local likelihood kernel estimator of degree $q=0$ (TTL0), degree $q=1$ (TTL1), and $q=2$ (TTL2), and $q=2$ with nearest neughbors (TTL2nn). We also considered the recent non-parametric approach by Mukhopadhyay and Parzen [Mukhopadhyay and Parzen, 2020] (MR), as well as the Probit transformation-based estimator from Geenens et al. [2017]. These approaches have R code on the CRAN (see Appendix E).
@@
To assess the performance of our model on synthetic bivariate datasets, we generated Gaussian copulas with correlations $\rho=0.1,0.5,0.9$, and Clayton and Frank copulas with parameters $\theta=1,5,10$. The marginal distributions were chosen as uncorrelated Normal distributions with parameters ($\mu=0,\sigma=1$). We generated 1,500 training samples and 500 test samples. Results are for test samples.
@@
The results are presented in Table 1, which contains the average negative log-likelihood on the test set for each copula. Results were estimated with a fixed seed of 30091985. The table also shows 95% Bootstrap Confidence Intervals [Wasserman, 2004]. Following suit with several guidelines, we choose to report CIs for better interpretability on the expected range of scores when replicating our studies [Gardner and Altman, 1986, Sterne and Smith, 2001, Altman et al., 2013, Ho et al., 2019, Dragicevic, 2016, Greenland et al., 2016, Imbens, 2021]. With 95% probability, a change of splits/seeds should achieve an average negative log-likelihood within the reported range.
@@
The winning 2-Cats models are highlighted in the table in Green. The best, on average, baseline methods are in Red. We consider a method better or worse than another when their confidence intervals do not overlap; thus, ties with the best model are shown in underscores. The table shows that 2-Cats is better than baseline methods when the dependency ($\rho$ or $\theta$) increases. For small dependencies, methods not based on Deep Learning outperform Deep ones (including 2-Cats). Nevertheless, 2-Cats is the winning method in 6 out of 9 datasets and is tied with the best in one other column.
5.3 2-Cats on Real Data
@@
Table 2: Negative log-likelihood on real datasets. $95\%$ confidence intervals (bootstrap) show. Green indicates that 2-Cats was, on average, better than the best baseline (indicated in Red). Underscores highlight ties with best model.
@@
As was done in previous work, we employ the Boston housing dataset, the Intel-Microsoft (INTC-MSFT), and Google-Facebook (GOOG-FB) stock ratios. These pairs are commonly used in Copula research; in particular, they are the same datasets used by [Ling et al., 2020, Ng et al., 2021]. We employed the same train/test and pre-processing as previous work for fair comparisons.
The results are presented in Table 2, again highlighting the superior performance of the 2-Cats models across all datasets. Only in one setting did the model underperform when compared with the literature. When we cross-reference results on synthetic and natural datasets, 2-Cats present’s itself as a promising novel approach to model Copulas.
@@ 6 Conclusions
In this paper, we presented Copula Approximating Transform models (2-Cats). Different from the literature when training our models; we focus not only on capturing the pseudo-likelihood of Copulas but also on several other properties, such as the partial derivatives of the $C$ function. We prove that our models are valid Copulas and thus useful for pairwise data modeling. Moreover, a second major innovation is proposing estimators that employ using Sobolev training. Overall, our results show the superiority of variations of 2-Cats on synthetic and real datasets.
@@
A natural follow-up for 2-Cats is on using the Pair Copula Construction (PCC) [Aas et al., 2009, Czado, 2010] to port Vines [Czado and Nagler, 2022, Low et al., 2018, Genest et al., 1995, Genest and Favre, 2007] to incorporate our method. PCC is a well-established algorithm to go from a 2D Copula to an ND One using Vines. Our focus on the main text of this paper was on showing 2-Cats as a valid 2D approach.
We also believe that evaluating other non-negative NNs for $m_{\bm{\theta}}$ is a promising direction [Marteau-Ferey et al., 2020, Rudi and Ciliberto, 2021, Tsuchida et al., 2023, Sladek et al., 2023, Loconte et al., 2023, Kortvelesy, 2023].
@@
We point out that before reaching our 2-Cats model, several variations of the model were considered. We refer the reader to Appendix D for results and details of such variations.
Appendix A Sampling from 2-Cats

### v4

@@ 5 Experimental Results
We now present our experimental results on different fronts: (1) validating our Empirical estimates for the first derivative, a crucial step when training 2-Cats; (2) evaluating 2-Cats on synthetic and real datasets without the lagrangian term; and, (3) evaluating the impact of the lagrangian term.
Before presenting results, we note that our models were implemented using Jax [^3] https://jax.readthedocs.io/en/latest//Flax [^4] https://flax.readthedocs.io. Given that Jax does not implement CDF methods for Multivariate Normal Distributions, we ported a fast approximation of bivariate CDFs [67] and employed it for our models that rely on Gaussian CDFs/Copulas. Our code is available at https://anonymous.4open.science/r/2cats-E765/. Baseline methods were evaluated using the author-supplied code (details are on Appendix G).
@@ 5.1 Empirical First Derivative Estimator
We validate our empirical estimations for the first derivative of Copulas. We use the closed-form equations for the first derivatives of Gaussian, Clayton, and Frank Copulas described in  [57].
These Copulas have a single parameter, which we varied. Gaussian Copulas ($\rho$) has a correlation parameter, and we set it to: 0.1, 0.5, and 0.9. Clayton and Frank copulas’ mixing parameter ($\theta$) was set to 1, 5, and 10. For each of these configurations, we measured the coefficient of determination ($R2$) between empirical estimates and exact formulae. Overall, we found that R2 was above $0.899$ in every setting, validating our approach.
5.2 2-Cats on Datasets (Without Lagrangian Term)
We now turn to our results on synthetic datasets. Our primary focus, as is the case on baseline methods, will be on capturing the data likelihood: $f(x_{1},x_{2})=k_{1}(x_{1})\,k_{2}(x_{2})\,\frac{\partial^{2}H_{\bm{\theta}}(u,v)}{\partial u\partial v}$ Here, $k_{1}(x_{1})$ and $k_{2}(x_{2})$ are KDE models for the PDFs for $X_{1}$ and $X_{2}$ respectivelly. The bandwidth for these methods is determined using Silverman’s rule [59].
As is commonly done, we evaluate the natural log of this quantity as the data log-likelihood and use its negative to compare 2-Cats with other methods. With this metric of interest in place, we considered two 2-Cats models: (1) 2-Cats-G (Final Layer Gaussian): 2-Cats with a Gaussian bivariate CDF as $G$; and, (2) 2-Cats-L (Final Layer Logistic): 2-Cats with a Logistic bivariate CDF as $G$.
Thus, we considered two forms of $G$. The first was the CDF of a bivariate Gaussian Distribution [67]. The second one is the CDF of the Flexible bivariate Logistic (see Section 11.8 of [4]):
$F(x_{1},x_{2})=\Pr[X_{1}<x_{1},X_{2}<x_{2}]=\big{(}(1+e^{-\alpha\frac{x_{1}-\mu_{1}}{\sigma_{1}}}+e^{-\alpha\frac{x_{2}-\mu_{2}}{\sigma_{2}}}+e^{-\alpha(\frac{x_{1}-\mu_{1}}{\sigma_{1}}+\frac{x_{2}-\mu_{2}}{\sigma_{2}})})^{\frac{1}{\alpha}}\big{)}^{-1}$
@@
Recall that the data density is proportional to the pseudo-likelihood. We primarily emphasized the likelihood loss ($w^{c}$). This accentuates our pivotal component. Notably, the scale of its values differs from that of the MSE components, necessitating a proportional adjustment in weight selection. In light of these factors, we fixed $w^{C}=0.01$, $w^{dC}=0.5$, and $w^{c}=0.1$. We note that these hyper-parameters were sufficient to show how our approach is better than the literature. Thus, no exhaustive grid search was needed. Our models were trained using early stopping. Here, for each training epoch, we evaluated the pseudo-log-likelihood on the training set and kept the weights of the epoch with the best likelihood on the training dataset.
Table 1: Results for Synthetic Data. Copula parameters are shown on the header. Marginals come from two Normal distributions ($\mu=0,\sigma=1$). $95\%$ confidence intervals (bootstrap) show. Green indicates that a 2-Cats model was, on average, better than the best baseline (indicated by Red). Underscores highlight the best statistical ties with the best model. We point out that data is normalized to the [0, 1] range before training. This is a necessity of some deep models such as [38, 51]. Non scaled comparisons are in Appendix E
@@
In both our synthetic and real data experiments, we considered the following approaches as Deep Learning baselines: Deep Archimedian Copulas (ACNET) [38]; Generative Archimedian Copulas (GEN-AC) [51]; Neural Likelihoods (NL) [13]; and Implicit Generative Copulas (IGC) [31]. Here, we use the same hyperparameters employed by the authors. We also considered several Non-Deep Learning baselines. These were the Parametric and Non-Parametric Copula Models from [49]. They are referenced as: The best Parametric approach from several known parametric copulas (Par), non-penalized Bernstein estimator (Bern), penalized Bernstein estimator (PBern), penalized linear B-spline estimator (PSPL1), penalized quadratic B-spline estimator (PSPL2), transformation local likelihood kernel estimator of degree $q=0$ (TTL0), degree $q=1$ (TTL1), and $q=2$ (TTL2), and $q=2$ with nearest neughbors (TTL2nn). We also considered the recent non-parametric approach by Mukhopadhyay and Parzen [47] (MR), as well as the Probit transformation-based estimator from [21].
@@
To assess the performance of our model on synthetic bivariate datasets, we generated Gaussian copulas with correlations $\rho=0.1,0.5,0.9$, and Clayton and Frank copulas with parameters $\theta=1,5,10$. The marginal distributions were chosen as uncorrelated Normal distributions with parameters ($\mu=0,\sigma=1$). We generated 1,500 training samples and 500 test samples. Results are for test samples. Some deep methods require datasets to be scaled to [0, 1] intervals [38, 51]. For this reason, we present results after a min-max scaling of the data.
@@
The results are presented in Table 1, which contains the average negative log-likelihood on the test set for each copula. Results were estimated with a fixed seed of 30091985. The table also shows 95% Bootstrap Confidence Intervals [70]. Following suit with several guidelines, we choose to report CIs for better interpretability on the expected range of [20, 65, 3, 28, 19, 24, 30].
@@
The winning 2-Cats models are highlighted in the table in Green. The best, on average, baseline methods are in Red. We consider a method better or worse than another when their confidence intervals do not overlap; thus, ties with the best model are shown in underscores. The table shows that 2-Cats is better than baseline methods when the dependency ($\rho$ or $\theta$) increases. For small dependencies, methods not based on Deep Learning outperform Deep ones (including 2-Cats). Nevertheless, 2-Cats is the winning method in 6 out of 9 datasets and is tied with the best in one case.
@@
Table 2: Negative log-likelihood on real datasets. $95\%$ confidence intervals (bootstrap) show. Green indicates that 2-Cats was, on average, better than the best baseline (indicated in Red). Underscores highlight ties with the best model. We point out that data is normalized to the [0, 1] range before training. This is a necessity of some deep models such as [38, 51]
@@
We now turn to real datasets. As was done in previous work, we employ the Boston housing dataset, the Intel-Microsoft (INTC-MSFT), and Google-Facebook (GOOG-FB) stock ratios. These pairs are commonly used in Copula research; in particular, they are the same datasets used by [38, 51]. We employed the same train/test and pre-processing as previous work for fair comparisons.
The results are presented in Table 2, again highlighting the superior performance of the 2-Cats models across all datasets. Only in one setting did the model underperform.
5.3 On the Lagrangian Term for P3
So far, 2-Cats presents itself as a promising approach for Copula modeling. Nevertheless, to sample from Copulas, we require that the marginals of the model come from a uniform distribution. This is our primary for the Lagrangian terms used to meet P3 (see Section 3).
We trained 2-Cats models on the three real-world datasets with and without the Lagrangian optimization of Eq (6). Due to space constraints, we present results for 2-Cats-L only. Here, models were trained for 1000 iterations. No early stopping was performed as our focus was on meeting P3.
We compare both the average absolute deviations (i.e., $abs_{u}=\frac{1}{n}\sum_{i}|H_{\bm{\theta}}(u_{i},1)-u_{i}|$ and $abs_{v}=\frac{1}{n}\sum_{i}|H_{\bm{\theta}}(1,v_{i})-v_{i}|$) as well as the average relative deviations (i.e., $rel_{u}=100\,\frac{1}{n}\sum_{i}|H_{\bm{\theta}}(u_{i},1)-u_{i}|/u_{i}$ and $rel_{v}=100\,\frac{1}{n}\sum_{i}|H_{\bm{\theta}}(1,v_{i})-v_{i}|/v_{i}$) for models trained with and without constraints.
For the GOOG-FB dataset, significant gains were achieved with the constraints, i.e.: $abs_{u}$ went from $0.17$ to $0.09$ and $abs_{v}$ improved from $0.16$ to $0.03$. $rel_{u}$ improved from $57\%$ to $35\%$ and $rel_{v}$ went from $54\%$ to $14\%$. The negative log-likelihood reduced slightly when Lagrangian was used (from -1.70 to -1.14). Nevertheless, the model is still better than the baselines (see Table 2). Results for the negative log-likelihood do not exactly match the previous ones, as no early stopping was done here.
For INTC-MSOFT, results were similar across $u$ and $v$ dimensions. As in both versions, no significant gains were observed in $abs_{u}$ nor $abs_{v}$; these scores are pretty small (below $0.009$) in both cases. In relative terms, an increase was observed in $abs_{v}$ from roughly 4% to 2%, with $abs_{u}$ being 4% in both cases. No significant changes were observed in the Boston dataset where errors were already minor $abs_{u}$ and $abs_{v}$ below $0.02$ in both instances, with relative errors ranging from 2% to 4% regardless of constraints being employed or not.
These results show that even 2-Cats without constraints may already approximate P3 (Boston and INTC-MSOFT). Constraints may present significant improvements in some cases as GOOG-FB.
@@ 6 Conclusions
In this paper, we presented Copula Approximating Transform models (2-Cats). Different from the literature when training our models; we focus not only on capturing the pseudo-likelihood of Copulas but also on meeting or approximating several other properties, such as the partial derivatives of the $C$ function. Moreover, a second major innovation is proposing estimators that employ using Sobolev training. Overall, our results show the superiority of 2-Cats on synthetic and real datasets.
@@
A natural follow-up for 2-Cats is on using the Pair Copula Construction (PCC) [1, 15] to port Vines [16, 41, 23, 22] to incorporate our method. PCC is a well-established algorithm to go from a 2D Copula to an ND One using Vines. We also believe that evaluating other non-negative NNs for $m_{\bm{\theta}}$ is a promising direction [45, 55, 68, 62, 40, 36].
@@
We point out that before reaching our 2-Cats model, several variations of the model were considered. We refer the reader to Appendix F for results and details of such variations.
Appendix A On Derivatives and Integrals of NNs
In this appendix, we present an example of the issue of approximating derivatives/integrals of NNs.
A.1 An Example on why approximating derivatives and integrals fail
We begin with a pictorial example of the issue. Before doing so, we take some time to revise the universal approximation properties of NNs [14, 29, 43]. We also shift our focus from Copulas for this argument. It is expected to state that the universal approximation theorem (UAT) guarantees that functions are approximated with NNs. However, given that when learning models from training data, it would be more correct to say that function evaluations at training points are approximated.
Figure 1: Models vs Derivatives. Here, we compare sample data from the simple $y=sin(x)$ process. Data is sampled in three settings, a “full” dataset with a larger range (left), a noisy dataset (middle), and a limited view (right). NN models (lines) approximate the data points in every case. However, the derivatives/integrals of the model do not approximate function derivatives (last row).
Consider a simple 4-Layer Relu NN with layers of 128 neurons. This will be our model. To avoid overfitting, this model will be optimized with a validation set. The problem with relying only on the UAT is shown in Figure 1. The figure shows a ground truth data-generating process of $y=\sin(x)+\eta$, where $\eta$ is some noise term. On the first column, when $\eta=0$, the model will approximate the data generating function (middle row), its derivative (bottom row), and integral (top row). When $\eta\sim N(0,0.5)$, the model will approximate the data points, but not necessarily the derivatives.
This example may seem to contradict [17, 29] others that looked into the universal approximation properties of NN regarding derivative. We point out that this is not the case. In Appendix A.3, we show how the variance will increase when estimating the derivative, leading to the figure’s oscillating behavior. From the figure, also note that integrals are approximated up to an asymptotic constant (see Appendix A.2, and  [39]); however, we still need to control for this constant (note the growing deviation on the top plot of the middle row).
We now present some results on the Integral (exemplified on the top row of Figure 1) and Derivative (bottom row) that tie in with our example. Furthermore, we briefly discuss the last column of the figure. Notice that on the extreme left and extreme right, the derivative of the network begins to diverge from the underlying process. The NN was not trained with points sampled after these boundaries. Under data shifts (e.g., data from a similar process not seen on the training set), the issues we point out throughout this Appendix should increase.
A.2 On the Integral of Neural Networks
To present a view of the theoretical bounds on the error of approximation of integrals, let $\hat{f}_{\bm{\theta}}(x)$ be some NN parametrized by $\bm{\theta}$. Also, let $f(x)$ represent our training dataset. We also make the reasonable assumption our dataset comes from some real function, $f^{\ast}(x)$, sampled under a symmetric additive noise, i.e., $f(x)=f^{\ast}(x)+\eta$, with an expected value equal to zero: $\mathbb{E}[\eta]=0$. Gaussian noise is such a case.
If a NN is a universal approximator [14, 29, 42, 10, 44] and it is trained on the dataset $f(x)$, we reach:
$\lVert\hat{f}_{\bm{\theta}}(x)-f(x)\rVert_{p}<\epsilon_{1}$
Most UAT proofs are valid for any $p$ norm ($p$ norms bound one another), and assuming the 1 norm, we reach: $\hat{f}_{\bm{\theta}}(x)=f(x)\pm\epsilon$ for some constant $\epsilon$. As a consequence, we reach:
$\hat{f}_{\bm{\theta}}(x)=f^{\ast}(x)+\eta\pm\epsilon_{1}$
Given that integration is a linear operator, we can show that:
$E_{\eta}[\int\hat{f}_{\bm{\theta}}(x)]=\int f^{\ast}(x)\pm\epsilon_{2}.$
$g(x)$ $=\int\hat{f}{{}_{\theta}}(x)\,dx$
[…]
$=\int E_{\eta}[f^{\ast}(x)]+E_{\eta}[\eta]\pm E_{\eta}[\epsilon_{1}]dx$
$=\int E_{\eta}[f^{\ast}(x)]\pm E_{\eta}[\epsilon_{1}]dx$
$=\int f^{\ast}(x)\pm\epsilon_{1}dx$
[…]
where $\epsilon_{2}$ is a constant as $\epsilon_{1}$ is also a constant. However, this constant will grow depending on the integrating interval. This growing error is visible on the plot on the center top of Figure 1.Controlling for $\epsilon_{2}$ will depend on the dataset’s quality and the training procedure.
A similar result is discussed in [39].
A.3 On the Derivative of Underlying Process and its impact on Neural Networks
To understand the oscillations on the derivatives of the example, let us rewrite the underlying process that generated the data as a Gaussian process. That is: $Y(x)=sin(x+\phi)+\eta$. $Y$ is the stochastic process and $\phi\sim Uniform(-2\pi,2\pi)$ is some phase component necessary for stationarity. The variance of this process will be given by the kernel function $K(\tau)$, which is now controlled by parameter $\tau$. Given that this process has the same variance over time, $x$, it is stationary, and a stationary Gaussian process has $K(\tau)=e^{-\frac{\tau}{2}}$ [37]. Before continuing, we note that in our example, we sampled from a Gaussian with a standard deviation of $\sigma=0.5$ and variance of $\sigma^{2}=0.25$. This leads to a value of $\tau$ that is of $\tau=\approx 2.8$ since $e^{-\frac{2.8}{2}}\approx 0.25$.
Now, let $Y^{\prime}(x)$ be the derivative of this process. This may be defined as:
$\frac{Y(x+h)-Y(x)}{h}$ $\rightarrow Y^{\prime}(x)$
if:
$E\big{[}(\frac{Y(x+h)-Y(x)}{h}-Y^{\prime}(x))^{2}\big{]}$ $\rightarrow 0$
when: $h\rightarrow 0$.
Given that our process is stationary and has the expected value of the noise term as zero, we can estimate the value of the derivative process as zero, i.e., $E[Y^{\prime}(x)]=0$ [37]. We also have that the variance is equal to the second derivative of $K(\tau)$: $Var[Y^{\prime}(x)]=\frac{\partial K}{\partial\tau\partial\tau}$ [37].
From here, we can derive $K$ twice to show that: $Var[Y^{\prime}(x)]=-e^{-\frac{\tau^{2}}{2}}+\tau^{2}e^{-\frac{\tau^{2}}{2}}$. Recall that $Var[Y(x)]=\sigma^{2}=0.25$ and $\tau\approx 2.8$. Now, by solving a simple inequality, we can show that the variance of the derivative process, $Var[Y^{\prime}(x)]$, is greater than the variance of the original process $Var[Y(x)]$, i.e., $Var[Y^{\prime}(x)]>Var[Y(x)]$ when $\tau>\sqrt{2}\approx 1.4$ (as is our example where $\tau\approx 2.8$).
This result explains the oscillation in the bottom middle plot of Figure 1. The derivative process has a higher variation; thus, we expect a higher variance in estimations of the derivative.
Overall, even if an NN is a universal approximation of derivatives [17, 29], the variance of the noise term will likely increase (here we began with a variance of 0.25 and it is already sufficient to show such an increase) even for the derivative of the NN.
Appendix B Sampling from 2-Cats